Non Constant Growth valuation

Hart Enterprises recently paid a dividend, D0, of $1.25. It expects to have non-constant growth of 20% for 2 years followed by a constant rate of 5% thereafter. The firm's required return is 8%. What is the firm's intrinsic value today?
Answer #1

Given for Hart Enterprises,

recent dividend P0 = $1.25

non-constant growth of 20% for 2 years

So, D1 = D0/1.2 = 1.25*1.2 = $1.5

D2 = D1*1.2 = $1.5*1.2 = $1.8

thereafter growth rate g = 5%

Firm's required return Ke = 8%

So, stock price at year 2 using constant dividend growth rate is

P2 = D2*(1+g)/(Ke-g) = 1.8*1.05/(0.08-0.05) = $63

intrinsic value of stock today is sum of PV of dividends and P2 discounted at Ke

So, intrinsic value P0 = D1/(1+Ke) + D2/(1+Ke)^2 + P2/(1+Ke)^2

P0 = 1.5/1.08 + 1.8/1.08^2 + 63/1.08^2 = $56.94

answered by: DoodleNyc